OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 55 miles per hour (mph), and the other car is traveling 40 mph. How long will it take before the cars are 285 miles apart? Answer 3 hrs 5 hrs 2 hrs 4 hrs

OpenStudy (sasogeek):

well 1 car is travelling at 55mph and the other at 40mph in the opposite direction, this implies that they are creating a space between themselves at a certain rate. obviously the rate is faster than their individual speeds because their moving away from each other so we can find that by adding their speeds, 55+40=95mph. we would have subtracted if they were moving in the same direction. now that we have the rate at which they're moving apart, we can find the time it would take them to become apart a certain distance. formula for time taken =distance/speed =>time taken to be apart 285miles = 285/95=3hours